A 20 kg box is pushed up a rough sloping ramp (30° relative to the incline) with a force of 300 N a, parallel to the horizontal with constant speed. Calculate the normal force on the block


1
Expert's answer
2021-06-07T14:16:11-0400


The net force on the y-direction is equal to zero, hence


"N-mg\cos 30^{\circ}-F\sin 30^{\circ}=0"

"N=mg\cos 30^{\circ}+F\sin 30^{\circ}"

"N=20\times 9.8\cos 30^{\circ}+300\sin 30^{\circ}=320\:\rm N"